OpenStudy (anonymous):

you have eight pens in your backpack. three are out of ink. if you select two pens at random, what is the probability that both have ink?

OpenStudy (dls):

So your n(S) = 8 No. of ink pens=3 P(1st ink pen)=3/8 P(2nd ink pen)=2/7 Required probability=\[\Large \frac{3}{8} \times \frac{2}{7}\]

OpenStudy (amoodarya):

\[p=\frac{ \left(\begin{matrix}5 \\ 2\end{matrix}\right) }{ \left(\begin{matrix}8 \\ 2\end{matrix}\right) }\]

OpenStudy (amoodarya):

answer is the same as "DLS" says

OpenStudy (anonymous):

=3/28

OpenStudy (goformit100):

A Warm Welcome to Open Study. Stay Studious.